# FedAPTA: Federated Multi-task Learning in Computing Power Networks with Adaptive Layer-wise Pruning and Task-aware Aggregation

### v1

@@ II-C Heterogeneous Model Recovery
where $\hat{w_{i}}$ is the model to be recovered and $w_{t}$ is the latest global model of task $t$, which is executed by device $i$. The term $(1-M_{i})$ effectively identifies the pruned parameters of local models, then the central server replaces these pruned values with their corresponding entries from the latest global model.
@@
II-D Task-aware Model Aggregation
In this section, we implement a task-aware aggregation method by device clustering, which is specifically designed to ensure the performance of the global model for multi-task scenarios in CPNs. In traditional FL settings, it is generally assumed that all devices are working on the same global task and hence can freely share updates for direct aggregation. However, in CPNs, participating devices are usually engaged in different tasks. This type of task heterogeneity presents a significant challenge for conventional aggregation methods, as directly combining models from different tasks may yield suboptimal or even misleading global models. Such models may fail to serve any individual task effectively, and in some cases, may even result in harmful interference across tasks. To solve this issue, the primary objective of our proposed method is to distinguish models that belong to different tasks. By doing so, the central server can perform task-aware aggregations, thereby preserving task-specific knowledge and avoiding negative transfer between unrelated tasks. It is particularly beneficial in multi-task FL scenarios in CPNs.
As aforementioned, for traditional FL frameworks like [3], which involves two fundamental processes: 1) The aggregation and broadcasting of models on the server side; 2) The training and uploading of models on the device side. Our modifications predominantly focus on the server side, while leaving the device-side operations unchanged. This design ensures that our method can be easily integrated into current FL frameworks in CPNs. Specifically, before the model aggregation step, we apply a model clustering step to group devices, thereby identifying which devices can collaboratively work on the same task.
$d_{cos}=1-\cos(\Delta w_{i},\Delta w_{j}),$ (4)
where $\Delta w_{i}=w_{i}-w_{t}$, and $\Delta w_{i},\Delta w_{j}$ are the local model updates on the device $i,j$ respectively. A small $d_{cos}$ value indicates that two devices share similar update directions and suggests a high probability of working on the same task, while a larger value indicates dissimilar updates and hence possible task divergence. Additionally, computing this distance across all model parameters can be computationally expensive and may include noisy signals from layers unrelated to task-specific features. According to pFedGraph [17], the last few fully connected layers are sufficient to reflect the fine-grained similarity of models, so we use the cosine distance of the parameters’ update values in the last few fully connected layers to approximate $d_{cos}$ to reduce computational cost. This approximation achieves a balance between clustering efficiency and performance by filtering out irrelevant noise from early layers.
After computing the cosine distances $d_{cos}$ between each pair of the received local models, a square matrix (i.e., distance matrix) is constructed where each entry represents the distance between a pair of models. Then the central server applies HDBSCAN [18] based on this matrix to identify groups of devices associated with the same task. Unlike traditional clustering algorithms, HDBSCAN operates without the number of clusters to be specified in advance, which aligns well with FL, where the number of distinct tasks may vary and is not known as a priori. Moreover, HDBSCAN is resilient to noise and outliers, making it particularly suitable for federated settings where the data may be uneven and noisy. Each resulting cluster ideally corresponds to a subset of devices working on the same task $t\in T$. The clustering process is described in Algorithm 3.
Subsequently, based on the clustering results, FedAPTA aggregates models for each task $t$ individually using a weighted average scheme. The weight assigned to each device is based on its task requirement and local dataset size, so that devices with larger dataset sizes play a more prominent role in updating the global model. The aggregation operation is formally defined as:
$w_{t}=\sum_{i\in C_{t}}\frac{|D_{i}|}{|D_{t}|}w_{i},$ (5)
III Experimental Evaluation
@@
In this section, we implement the proposed FedAPTA framework on a real-world FL platform, FedML [19], followed by a performance comparison with nine existing frameworks. Then we investigate how varying pruning ratios affect the learning performance of FedAPTA. Finally, we evaluate the effectiveness of various model similarity metrics in distinguishing models belonging to different tasks.
III-A Experimental Setting
III-A1 FL environment
@@
We simulate fifty heterogeneous devices and set up five distinct classification tasks, with each task being assigned to a group of ten devices. We employ ResNet18 [20] and ShuffleNetV2 [21] models to carry out these classification tasks. Following the setup of FedLPS [13], we use the pretrained model on ImageNet [22] as the initialized global model and freeze the first quarter of the layers during training. Moreover, inspired by FedDC [23], we replace the Batch Normalization layers in the model with Group Normalization [24] layers, and learn a local drift variable to track and correct the gap between the local model and the global model, in order to mitigate the effect of the local data distributions on model performance. In the practical application under CPNs, devices participating in FL can be selected based on their available computing resources collected. Therefore, in our simulation experiments, the FL training process includes all heterogeneous devices. Experimental evaluations are performed on a GPU server equipped with four NVIDIA RTX 3090 GPUs.
TABLE II: Hyper-parameters used in FL training.
@@
TABLE III: The ResNet18 model accuracy (%) comparison under both i.i.d. and non-i.i.d. settings.
@@
i.i.d. partition FedAvg [3] 92.85 87.72 79.04 70.79 79.29 81.94
Ditto [25] 93.06 87.73 79.21 70.66 79.30 82.02
FedProx [26] 93.01 87.69 79.29 70.52 79.26 81.96
FedGen [27] 92.94 87.86 79.27 70.37 79.18 81.93
MOON [28] 92.93 87.63 79.32 70.43 79.32 81.93
FedBABU [29] 92.88 87.54 79.25 70.45 79.37 81.90
FedNTD [30] 93.07 87.55 79.16 70.66 79.21 81.93
FedLC [31] 92.97 87.63 79.21 70.66 79.29 81.95
FedLPS [13] 98.06 89.18 91.64 80.10 85.01 88.80
@@
Non-i.i.d. partition FedAvg [3] 91.73 86.46 76.02 65.35 77.2 79.35
Ditto [25] 91.82 86.38 75.89 65.98 77.19 79.45
FedProx [26] 91.73 86.44 76.04 65.82 77.2 79.45
FedGen [27] 91.9 86.43 75.99 65.95 77.03 79.46
MOON [28] 91.75 86.5 75.92 65.64 77.23 79.4
FedBABU [29] 91.82 86.62 76.15 66.35 77.5 79.69
FedNTD [30] 91.65 86.5 76.14 65.55 77.18 79.4
FedLC [31] 91.76 86.26 76 65.81 77.09 79.38
FedLPS [13] 97.25 84.24 88.22 67.49 79.65 83.37
@@
TABLE IV: The ShuffleNetV2 model accuracy (%) comparison under both i.i.d. and non-i.i.d. settings.
@@
i.i.d. partition FedAvg [3] 88.21 82.62 66.76 58.45 75.48 74.30
Ditto [25] 88.07 82.61 67.38 57.99 75.39 74.29
FedProx [26] 88.37 82.59 67.64 58.82 75.38 74.56
FedGen [27] 88.15 82.65 67.25 58.15 75.34 74.31
MOON [28] 88.27 82.61 67.80 57.91 75.50 74.42
FedBABU [29] 88.16 82.89 67.89 57.90 75.75 74.52
FedNTD [30] 87.74 82.71 67.53 57.63 75.47 74.22
FedLC [31] 88.14 82.68 67.60 57.76 75.43 74.32
FedLPS [13] 84.57 79.86 71.99 66.03 83.50 77.19
@@
Non-i.i.d. partition FedAvg [3] 85.13 80.25 63.06 52.57 72.65 70.73
Ditto [25] 85.44 80.06 63.18 53.49 72.76 70.99
FedProx [26] 85.84 79.92 63.20 52.58 72.80 70.87
FedGen [27] 85.40 79.91 62.93 52.56 72.83 70.72
MOON [28] 85.82 79.86 63.24 52.40 72.78 70.82
FedBABU [29] 85.35 80.34 62.92 53.54 73.21 71.07
FedNTD [30] 85.48 80.07 62.98 52.13 72.76 70.69
FedLC [31] 85.31 80.31 63.62 52.61 72.81 70.93
FedLPS [13] 78.49 74.07 65.09 59.71 77.55 70.98
@@
III-A2 Datasets and data partition
We explore on five widely recognized datasets: MNIST [32], FashionMNIST [33], SVHN [34], CIFAR10 [35], and EMNIST [36] to simulate the classification tasks. Under the independent and identically distributed (i.i.d.) setting, each dataset was evenly allocated across the devices. In the non-independent and identically distributed (non-i.i.d.) setting, following [13, 37], we employ the Latent Dirichlet Allocation (LDA) method to construct the non-i.i.d. data. Within LDA, we use the conventional setting of $\boldsymbol{\alpha}=0.5$ to simulate non-i.i.d. data distributions, where $\alpha$ is utilized to regulate the level of data heterogeneity.
III-A3 Comparison frameworks
We compare the proposed FedAPTA framework with FedAvg [3], Ditto [25], FedProx [26], FedGen [27], MOON [28], FedBABU [29], FedNTD [30], FedLC [31] and FedLPS [13]. FedAvg is a classical FL framework that requires each device to update the whole model, which consequently results in computing resources wastage within multi-task FL environments in CPNs. Ditto enables devices to maintain personalized models, thereby adapting effectively to the local data distributions. FedProx introduces a regularization term to encourage local models to stay close to the global model from the same task, while allowing devices with limited computational capabilities to perform fewer local updates. FedGen addresses data heterogeneity issues in FL by training a generator on the server side that aggregates knowledge from devices as prior knowledge for local training of the same task. MOON employs contrastive learning on model representations, utilizing the discrepancies between the current local model and the global model from the same task to guide local training, effectively mitigating challenges arising from non-i.i.d. data distributions. FedBABU incorporates attention-based representation learning to enhance the model’s representational power, thereby improving performance in image classification tasks. FedNTD mitigates the forgetting problem of the global model during the FL process by preserving non-ground-truth class information during local training, which leads to improved performance on non-i.i.d. data. FedLC calibrates logits to reduce model overfitting on minority classes within a task, enhancing classification accuracy under long-tail or imbalanced data distributions. FedLPS reduces resource consumption for multi-task learning in heterogeneous FL through local parameter sharing and a channel-wise model pruning algorithm. To ensure reliability, experiments are repeated three times to derive average results. Moreover, to ensure a just comparison, the same initialized global model was adopted for both FedAPTA and all comparative frameworks during the experiments. Detailed information regarding the training hyperparameters is provided in Table II.
@@ III-A3 Comparison frameworks
Figure 2: Model accuracy of FedAPTA, FedLPS [13], FedDrop [9], and FedRolex [10] on the RseNet18 model with different pruning ratios on the non-i.i.d. setting of the SVHN, CIFAR10, and EMNIST datasets.
(a) Non-i.i.d.
(b) i.i.d.
TABLE V: Number of parameters of ResNet18 and ShuffleNetV2 under different pruning ratios in FedAPTA.
@@
III-B Comparison of the Model Accuracy
[…]
III-C Effect of Pruning Ratios
In FL frameworks where model pruning is used, the number of parameters in the model to be trained is significantly reduced, thereby decreasing the computing resources required for training local models, as demonstrated in Table V. However, as the pruning ratio increases, the model’s accuracy tends to deteriorate. Therefore, in this subsection, we investigate how varying pruning ratios affect model performance. In this experiment, we compare the proposed adaptive layer-wise model pruning method to the pruning methods used in FedLPS [13], FedDrop [9], and FedRolex [10], under the setting of $\rho=\{0.2,0.4,0.6,0.8\}$. Fig. 2 illustrates the model accuracy of the ResNet-18 model after 100 communication rounds of FL under non-i.i.d. settings on the SVHN, CIFAR-10, and EMNIST datasets. The results indicate that, on these three datasets, FedAPTA consistently outperforms the aforementioned pruning methods when the pruning ratio $\rho$ takes values between 0.2 and 0.8, and is capable of maintaining satisfactory model accuracy even under relatively high pruning rates.
III-D Effects of Model Similarity Metrics

### v2

@@ 2.3 Heterogeneous Model Recovery
where $\hat{w_{i}}$ is the model to be recovered and $w_{t}$ is the latest global model of task $t$, which is executed by device ii. The term $(1-M_{i})$ effectively identifies the pruned parameters of local models, then the central server replaces these pruned values with their corresponding entries from the latest global model.
@@
2.4 Task-aware Model Aggregation
In this section, we implement a task-aware aggregation method by device clustering, which is specifically designed to ensure the performance of the global model for multi-task scenarios. In traditional FL settings, it is generally assumed that all devices are working on the same global task and hence can freely share updates for direct aggregation. However, in real-world scenarios, participating devices are usually engaged in different tasks. This type of task heterogeneity presents a significant challenge for conventional aggregation methods, as directly combining models from different tasks may yield suboptimal or even misleading global models. Such models may fail to serve any individual task effectively, and in some cases, may even result in harmful interference across tasks. To solve this issue, the primary objective of our proposed method is to distinguish models that belong to different tasks. By doing so, the central server can perform task-aware aggregations, thereby preserving task-specific knowledge and avoiding negative transfer between unrelated tasks. It is particularly beneficial in multi-task FL scenarios.
As aforementioned, for traditional FL frameworks like [mcmahan2017communicationfedavg], which involves two fundamental processes: 1) The aggregation and broadcasting of models on the server side; 2) The training and uploading of models on the device side. Our modifications predominantly focus on the server side, while leaving the device-side operations unchanged. This design ensures that our method can be easily integrated into current FL frameworks. Specifically, before the model aggregation step, we apply a model clustering step to group devices, thereby identifying which devices can collaboratively work on the same task.
$d_{cos}=1-\cos(\Delta w_{i},\Delta w_{j}),$ (4)
where $\Delta w_{i}=w_{i}-w_{t}$, and $\Delta w_{i},\Delta w_{j}$ are the local model updates on the device $i,j$ respectively. A small $d_{cos}$ value indicates that two devices share similar update directions and suggests a high probability of working on the same task, while a larger value indicates dissimilar updates and hence possible task divergence. Additionally, computing this distance across all model parameters can be computationally expensive and may include noisy signals from layers unrelated to task-specific features. According to pFedGraph [ye2023personalized_pfedgraph], the last few fully connected layers are sufficient to reflect the fine-grained similarity of models, so we use the cosine distance of the parameters’ update values in the last few fully connected layers to approximate $d_{cos}$ to reduce computational cost. This approximation achieves a balance between clustering efficiency and performance by filtering out irrelevant noise from early layers.
After computing the cosine distances $d_{cos}$ between each pair of the received local models, a square matrix (i.e., distance matrix) is constructed where each entry represents the distance between a pair of models. Then the central server applies HDBSCAN [campello2013density_hdbscan] based on this matrix to identify groups of devices associated with the same task. Unlike traditional clustering algorithms, HDBSCAN operates without the number of clusters to be specified in advance, which aligns well with FL, where the number of distinct tasks may vary and is not known as a priori. Moreover, HDBSCAN is resilient to noise and outliers, making it particularly suitable for federated settings where the data may be uneven and noisy. Each resulting cluster ideally corresponds to a subset of devices working on the same task $t\in T$. The clustering process is described in Algorithm 3.
Subsequently, based on the clustering results, FedAPTA aggregates models for each task $t$ individually using a weighted average scheme. The weight assigned to each device is based on its local dataset size, so that devices with larger dataset sizes play a more prominent role in updating the global model. The aggregation operation is formally defined as:
$w_{t}=\sum_{i\in C_{t}}\frac{|D_{i}|}{|D_{t}|}w_{i},$ (5)
@@ 2.4 Task-aware Model Aggregation
3 Experimental Evaluation
@@
In this section, we implement the proposed FedAPTA framework on a real-world FL platform, FedML [he2020fedml], followed by a performance comparison with nine existing frameworks. Then we investigate how varying pruning ratios affect the learning performance of FedAPTA. Finally, we evaluate the effectiveness of various model similarity metrics in distinguishing models belonging to different tasks.
3.1 Experimental Setting
3.1.1 FL environment
@@
We simulate fifty heterogeneous devices and set up five distinct classification tasks, with each task being assigned to a group of ten devices. We employ ResNet18 [he2016deepresnet] and ShuffleNetV2 [zhang2018shufflenet] models to carry out these classification tasks. Following the setup of FedLPS [jia2024fedlps], we use the pretrained model on ImageNet [russakovsky2015imagenet] as the initialized global model and freeze the first quarter of the layers during training. Moreover, inspired by FedDC [gao2022feddc], we replace the Batch Normalization layers in the model with Group Normalization [wu2018group] layers, and learn a local drift variable to track and correct the gap between the local model and the global model, in order to mitigate the effect of the local data distributions on model performance. In our simulation experiments, the FL training process includes all heterogeneous devices. Experimental evaluations are performed on a GPU server equipped with four NVIDIA RTX 3090 GPUs.
Table 2: Hyper-parameters used in FL training.
@@
3.1.2 Datasets and data partition
We explore on five widely recognized datasets: MNIST [lecun1998gradientmnist], FashionMNIST [xiao2017fashionmnist], SVHN [netzer2011readingsvhn], CIFAR10 [krizhevsky2009learningcifar10], and EMNIST [cohen2017emnist] to simulate the classification tasks. Under the independent and identically distributed (i.i.d.) setting, each dataset was evenly allocated across the devices. In the non-independent and identically distributed (non-i.i.d.) setting, following [jia2024fedlps, luo2021noDirichlet], we employ the Latent Dirichlet Allocation (LDA) method to construct the non-i.i.d. data. Within LDA, we use the conventional setting of $\boldsymbol{\alpha}=0.5$ to simulate non-i.i.d. data distributions, where $\alpha$ is utilized to regulate the level of data heterogeneity.
3.1.3 Comparison frameworks
We compare the proposed FedAPTA framework with FedAvg [mcmahan2017communicationfedavg], Ditto [li2021ditto], FedProx [li2020federatedfedprox], FedGen [zhu2021datafedgen], MOON [li2021modelmoon], FedBABU [oh2022fedbabu], FedNTD [lee2022preservation_fedntd], FedLC  [zhang2022federated_fedlc] and FedLPS [jia2024fedlps]. FedAvg is a classical FL framework that requires each client to update the entire model, thus the training efficiency of FL with heterogeneous devices is affected by participants with lower computational capabilities. Ditto enables devices to maintain personalized models, thereby adapting effectively to the local data distributions. FedProx introduces a regularization term to encourage local models to stay close to the global model from the same task, while allowing devices with limited computational capabilities to perform fewer local updates. FedGen addresses data heterogeneity issues in FL by training a generator on the server side that aggregates knowledge from devices as prior knowledge for local training of the same task. MOON employs contrastive learning on model representations, utilizing the discrepancies between the current local model and the global model from the same task to guide local training, effectively mitigating challenges arising from non-i.i.d. data distributions. FedBABU incorporates attention-based representation learning to enhance the model’s representational power, thereby improving performance in image classification tasks. FedNTD mitigates the forgetting problem of the global model during the FL process by preserving non-ground-truth class information during local training, which leads to improved performance on non-i.i.d. data. FedLC calibrates logits to reduce model overfitting on minority classes within a task, enhancing classification accuracy under long-tail or imbalanced data distributions. FedLPS reduces resource consumption for multi-task learning in heterogeneous FL through local parameter sharing and a channel-wise model pruning algorithm. To ensure reliability, experiments are repeated three times to derive average results. Moreover, to ensure a just comparison, the same initialized global model was adopted for both FedAPTA and all comparative frameworks during the experiments. Detailed information regarding the training hyperparameters is provided in Table 2.
Table 3: The ResNet18 model accuracy (%) comparison under both i.i.d. and non-i.i.d. settings.
@@ 3.1.3 Comparison frameworks
i.i.d. partition FedAvg [mcmahan2017communicationfedavg] 92.85 87.72 79.04 70.79 79.29 81.94
Ditto [li2021ditto] 93.06 87.73 79.21 70.66 79.30 82.02
FedProx [li2020federatedfedprox] 93.01 87.69 79.29 70.52 79.26 81.96
FedGen [zhu2021datafedgen] 92.94 87.86 79.27 70.37 79.18 81.93
MOON [li2021modelmoon] 92.93 87.63 79.32 70.43 79.32 81.93
FedBABU [oh2022fedbabu] 92.88 87.54 79.25 70.45 79.37 81.90
FedNTD [lee2022preservation_fedntd] 93.07 87.55 79.16 70.66 79.21 81.93
FedLC [zhang2022federated_fedlc] 92.97 87.63 79.21 70.66 79.29 81.95
FedLPS [jia2024fedlps] 98.06 89.18 91.64 80.10 85.01 88.80
@@
Non-i.i.d. partition FedAvg [mcmahan2017communicationfedavg] 91.73 86.46 76.02 65.35 77.2 79.35
Ditto [li2021ditto] 91.82 86.38 75.89 65.98 77.19 79.45
FedProx [li2020federatedfedprox] 91.73 86.44 76.04 65.82 77.2 79.45
FedGen [zhu2021datafedgen] 91.9 86.43 75.99 65.95 77.03 79.46
MOON [li2021modelmoon] 91.75 86.5 75.92 65.64 77.23 79.4
FedBABU [oh2022fedbabu] 91.82 86.62 76.15 66.35 77.5 79.69
FedNTD [lee2022preservation_fedntd] 91.65 86.5 76.14 65.55 77.18 79.4
FedLC [zhang2022federated_fedlc] 91.76 86.26 76 65.81 77.09 79.38
FedLPS [jia2024fedlps] 97.25 84.24 88.22 67.49 79.65 83.37
@@
3.2 Comparison of the Model Accuracy
We compare the model accuracy of FedAPTA with FedAvg, Ditto, FedProx, FedGen, MOON, FedBABU, FedNTD, FedLC, and FedLPS under both i.i.d. and non-i.i.d. settings. In this experiment, pruning ratios $\rho$ of $\{0,0.2,0.4,0.6,0.8\}$ were assigned to every group of five devices. Table 3 presents the accuracy results of the ResNet18 model on i.i.d. and non-i.i.d. datasets, while Table 4 details the accuracy outcomes for the ShuffleNetV2 model. The superiority of FedAPTA can be attributed to three key factors: first, the models tailored for devices undergo meticulous pruning to preserve their performance as much as possible. Second, the heterogeneous model recovery algorithm uses information contained within the global model to assist in aggregating local models, facilitating improved knowledge transfer, and enabling models to learn more effectively from other devices. Third, the central server computes the cosine similarity among all received local models and employs HDBSCAN based on these similarities to cluster local models belonging to different tasks separately, thereby guaranteeing that the aggregated global models can correctly learn from knowledge based on their assigned tasks.
Table 4: The ShuffleNetV2 model accuracy (%) comparison under both i.i.d. and non-i.i.d. settings.
@@ 3.2 Comparison of the Model Accuracy
i.i.d. partition FedAvg [mcmahan2017communicationfedavg] 88.21 82.62 66.76 58.45 75.48 74.30
Ditto [li2021ditto] 88.07 82.61 67.38 57.99 75.39 74.29
FedProx [li2020federatedfedprox] 88.37 82.59 67.64 58.82 75.38 74.56
FedGen [zhu2021datafedgen] 88.15 82.65 67.25 58.15 75.34 74.31
MOON [li2021modelmoon] 88.27 82.61 67.80 57.91 75.50 74.42
FedBABU [oh2022fedbabu] 88.16 82.89 67.89 57.90 75.75 74.52
FedNTD [lee2022preservation_fedntd] 87.74 82.71 67.53 57.63 75.47 74.22
FedLC [zhang2022federated_fedlc] 88.14 82.68 67.60 57.76 75.43 74.32
FedLPS [jia2024fedlps] 84.57 79.86 71.99 66.03 83.50 77.19
@@
Non-i.i.d. partition FedAvg [mcmahan2017communicationfedavg] 85.13 80.25 63.06 52.57 72.65 70.73
Ditto [li2021ditto] 85.44 80.06 63.18 53.49 72.76 70.99
FedProx [li2020federatedfedprox] 85.84 79.92 63.20 52.58 72.80 70.87
FedGen [zhu2021datafedgen] 85.40 79.91 62.93 52.56 72.83 70.72
MOON [li2021modelmoon] 85.82 79.86 63.24 52.40 72.78 70.82
FedBABU [oh2022fedbabu] 85.35 80.34 62.92 53.54 73.21 71.07
FedNTD [lee2022preservation_fedntd] 85.48 80.07 62.98 52.13 72.76 70.69
FedLC [zhang2022federated_fedlc] 85.31 80.31 63.62 52.61 72.81 70.93
FedLPS [jia2024fedlps] 78.49 74.07 65.09 59.71 77.55 70.98
@@
Figure 2: Model accuracy of FedAPTA, FedLPS [jia2024fedlps], FedDrop [caldas2018expanding_feddrop], and FedRolex [alam2022fedrolex] on the RseNet18 model with different pruning ratios on the non-i.i.d. setting of the SVHN, CIFAR10, and EMNIST datasets.
Table 5: Number of parameters of ResNet18 and ShuffleNetV2 under different pruning ratios in FedAPTA.
@@
3.3 Effect of Pruning Ratios
In FL frameworks where model pruning is used, the number of parameters in the model to be trained is significantly reduced, thereby decreasing the computing resources required for training local models, as demonstrated in Table 5. However, as the pruning ratio increases, the model’s accuracy tends to deteriorate. Therefore, in this subsection, we investigate how varying pruning ratios affect model performance. In this experiment, we compare the proposed adaptive layer-wise model pruning method to the pruning methods used in FedLPS [jia2024fedlps], FedDrop [caldas2018expanding_feddrop], and FedRolex [alam2022fedrolex], under the setting of $\rho=\{0.2,0.4,0.6,0.8\}$. Fig. 2 illustrates the model accuracy of the ResNet-18 model after 100 communication rounds of FL under non-i.i.d. settings on the SVHN, CIFAR-10, and EMNIST datasets. The results indicate that, on these three datasets, FedAPTA consistently outperforms the aforementioned pruning methods when the pruning ratio $\rho$ takes values between 0.2 and 0.8, and is capable of maintaining satisfactory model accuracy even under relatively high pruning rates.
3.4 Effects of Model Similarity Metrics